# Scatterless interferences: Delay of laminar-to-turbulent flow transition by a lattice of subsurface phonons

### v2

@@ Abstract
Turbulent boundary-layer flows generate significantly higher surface-friction drag than flows that remain laminar. Avoidance or delay of laminar-to-turbulent transition is therefore critical for the efficient operation of air and water vehicles. However, control of the flow perturbations that trigger this transition is difficult due to their unstable and persistent nature. Here we present a passive, phonons-based approach that targets these perturbations through engineered interactions at the flow-surface interface. This intervention is achieved via a tailored frequency- and wavevector-dependent coupling between the flow and a smooth, homogeneous surface, beneath which a functional material architecture is entirely embedded. The subsurface consists of a wall-parallel lattice of wall-normal phononic structural units. Each discrete phononic subsurface (PSub) unit is designed to respond to local flow disturbances in an out-of-phase manner, effectively suppressing the perturbations at the point of interaction. Collectively, the lattice induces interference effects that cause the kinetic energy of the flow instabilities to decay downstream of the intervention region, and thereby delaying the transition. To guide the design of the subsurface lattice, a Bloch-wave analysis is developed for the flow perturbations within a representative unit cell interfaced with a PSub unit. This is followed by direct numerical simulations to validate the concept. This approach represents a paradigm shift in the design philosophy for aircraft, ship, and submarine surfaces--moving beyond streamlined shaping to leveraging subsurface phonon engineering for the reduction of drag and enhancement of overall performance.
Significance
Airflow over aircraft wings or around the fuselage generates skin-friction drag, severely reducing fuel efficiency. Similar drag effects occur in water flow around ships and submarines and in the transport of gas or oil through long-range pipelines. This friction drag can be significantly reduced by maintaining laminar flow or delaying the transition to turbulence. In this work, we demonstrate that phonon motion—low-amplitude intrinsic vibrations—within a lattice of architected materials embedded just beneath the flow-exposed surface can be engineered to fundamentally alter the fluid-structure interaction. This, in turn, enables the suppression of the flow instabilities responsible for triggering the transition to turbulence. The phonon properties of each subsurface unit in the lattice are precisely tailored to influence the flow locally, while the collective arrangement is engineered to produce a favorable downstream interference effect within the flow. This approach may alternatively be adapted for flow destabilization to delay separation or to enhance chemical mixing and combustion, offering a versatile platform for a broad range of flow control applications.
@@ Introduction
The interaction between a flowing fluid and a surface is a dynamical process central to a broad range of both natural and artificial systems. Two most consequential examples are flow over an aircraft wing and around a ship’s hull. The nature of the fluid-structure interaction in these cases governs critical phenomena such as the intensity of skin-friction drag, the onset location of laminar-to-turbulent transition, and the likelihood of flow separation—all of which have significant implications for overall vehicle performance [1, 2]. Skin-friction drag, in particular, carries substantial economic and environmental costs, as it directly impacts fuel efficiency. Effective intervention that delays flow transition is thus vital because wall-bounded turbulent flows may generate an order-of-magnitude or more larger surface-friction drag compared to laminar flows. On the other hand, flow separation influences aerodynamic and hydrodynamic stability, form drag, and vehicle maneuverability. Targeted intervention is therefore highly beneficial for control of flow separation in these applications. Similar fluid-structure-interaction challenges are encountered in various other contexts, including land vehicles, wind turbines, long-range pipelines, processes requiring enhanced turbulent mixing, among others.
@@
In this work, we present the concept of a lattice of PSubs comprising a collection of individual PSub units laid out following a square or hexagonal lattice symmetry  [^3] Other lattice symmetries may also be considered in the future.. Similar to acoustic or elastic waves propagating around or through a lattice of rigid or elastic scatterers, respectively (see Fig. 1a), we design our configuration on the basis of a rigorous Bloch wave analysis where the field variable is the flow perturbations. Yet, in our system there are no scatterers. In contrast to classical scattering problems where the interfering waves and the scattering objects are in the same space, the flow instability waves interfere as a result of the local influence of each PSub on the perturbation velocity components. Here we recall that the PSubs are located beneath the spatial domain of the flow (see Fig. 1b). After developing Bloch’s theorem for this unique problem, we demonstrate by direct numerical simulation (DNS) of the coupled fluid-PSub lattice systems sustained downstream reduction in the perturbation kinetic energy (KE), which indicates a delay in laminar-to-turbulent transition. Moreover, as dictated by the PSub lattice symmetry, this approach is immune to any changes in the direction of propagation of the instability waves and may be tuned to accommodate a relatively wide range of wavenumbers along each direction--similar to the classical acoustic Bragg scattering problem which we also demonstrate to provide a direct analogy and comparison.
@@ Results
We form our lattice of PSubs using identical PSub units, all installed in a channel model. We select a channel flow problem for simplicity, but all the underlying concepts are readily applicable to boundary layer flows. The design of the PSub unit used in this investigation is shown in Fig. 2a. It consists of an elastic rod, composed of five unit cells, with each unit cell accommodating a local resonator (which in practice may be realized as a cantilevered pillar).The dispersion curves and the amplitude and phase frequency-response characteristics for this nominal PSub configuration is shown in Fig. 2b. The product of the amplitude and phase yields a performance metric [23] which provides an a priori prediction of the behavior of the PSub once passively engaged with the flow--this quantity is also plotted in Fig. 2b.
@@ Discussion
In conclusion, we have shown that by engineering the phonon properties of the subsurface--virtually entirely concealed beneath the flow-exposed surface--it is possible to exert full command over the very nature of the fluid-structure interaction. This passive intervention offers precise mechanistic and tunable control of the behavior of wall-bounded flows. Our findings demonstrate that the principles of Bloch wave analysis, traditionally applied to electronic, photonic, acoustic, and elastic waves [32, 33], can be effectively extended to flow perturbations influenced by PSubs. By leveraging locally resonant metamaterial-based individual PSub units, we have shown that flow instabilities can be manipulated through interference effects, in a manner analogous to wave phenomena in periodic elastic or acoustic media. This approach enables passive control over critical processes such as the transition to turbulence, with the PSub lattice facilitating attenuation of unstable T-S waves downstream, while being robust to variations in wavenumber and direction of propagation. Other types of flow instabilities may be similarly treated in future studies. With the ongoing influx of new concepts from phonon engineering and metamaterials [34, 35, 36], the prospects of further future improvements of the performance following this framework is highly accessible, and so is the path to practical implementation given the rapid advancements in fabrication technology [37, 38]. A wide range of flow-control applications stand to benefit, spanning from hypersonic airflows to low-speed liquid transport. It is conceivable that within this framework future aircrafts will maintain a fully laminar boundary layer over their surfaces, enabling transformative improvements in fuel efficiency. In addition, PSub lattices may be simultaneously designed to inhibit aeroacoustic emissions, offering the additional rewards of improved passenger comfort and reduced environmental noise pollution. Similar advances will materialize for hydroacoustic emissions as well.

### v3

# Scatterless interference: Delay of laminar-to-turbulent flow transition by a lattice of subsurface phonons

@@ Abstract
Wave interference has historically relied on scattering objects placed within the wave domain. Here, we introduce a fundamentally new mechanism: scatterless interference induced by a lattice of subsurface phonon motion beneath a smooth wall interfacing with a transitioning boundary-layer flow. The subsurface consists of a wall-parallel lattice of wall-normal frequency-dependent phononic structural units, each designed to respond to local flow perturbations in an out-of-phase manner, suppressing them at the point of interaction. Collectively, the lattice induces interference effects that cause the kinetic energy of flow instabilities to decay downstream, thereby delaying laminar-to-turbulent transition. To guide the design of the phononic subsurface lattice, a Bloch-wave unit-cell analysis is developed for the flow perturbations, and direct numerical simulations validate the concept. This work establishes scatterless interference as a distinct physical phenomenon and represents a paradigm shift in the design of aerodynamic and hydrodynamic surfaces—moving beyond streamlined shaping to leveraging subsurface phonon engineering for drag reduction and enhanced performance.
@@ Introduction
The control of wave interference through the deliberate placement of scattering objects has been a cornerstone principle in wave physics for over two centuries. Its classical origins trace back to Thomas Young’s double-slit experiment [1], which first demonstrated the wave nature of light via interference between scattered paths. This concept evolved into a general principle applied across multiple disciplines. In electromagnetics, the development of photonic crystals [2, 3] relied on arrays of dielectric scatterers directly interacting with the wave field to trigger spatial constructive and destructive interferences that bring rise to Bragg scattering. A similar mechanism emerged in acoustics and elasticity, where periodic voids or inclusions, forming phononic crystals [4, 5], have enabled the formation of wave interference patterns and spectral band gaps through in-domain scattering. Even in quantum mechanics, interference effects typically arise through the interaction of wavefunctions with potentials or obstacles embedded in the same domain of motion. These classical and modern mechanisms share a foundational attribute: the structures responsible for interference are located within the same spatial domain as the wave field they manipulate  [^1] The only exception is having active external excitations to trigger the scattering [55, 56, 57]. To date, this paradigm has remained essentially unchallenged across physics. In this work, we introduce a fundamentally new mechanism of passive wave interference that breaks from the conventional paradigm: scatterless interference induced by a lattice of phononic subsurface units. In sharp contrast to prevalent scattering mechanisms, the interference patterns in our fully passive system arise not from scatterers embedded within the wave domain, but from elastic phononic structures buried beneath the surface—outside the spatial domain occupied by the propagating waves. We present this new physical mechanism in the context of the long-standing classical fluid dynamics problem of laminar-to-turbulent transition.
Fluid-structure interaction
@@
Scatterless interference
In this work, we present the concept of a lattice of PSubs comprising a collection of individual PSub units laid out following a square or hexagonal lattice symmetry  [^4] Other lattice symmetries may also be considered in the future.. Similar to acoustic or elastic waves propagating around or through a lattice of rigid or elastic scatterers, respectively (see Fig. 1a), we design our configuration on the basis of a rigorous Bloch wave analysis where the field variable is the flow perturbations. Yet, in our system there are no scatterers. In contrast to classical scattering problems where the interfering waves and the scattering objects are located in the same space, the flow instability waves interfere as a result of the local influence of each PSub on the perturbation velocity components. Here we recall that the PSubs are located beneath the spatial domain of the flow (see Fig. 1b). After developing Bloch’s theorem for this unique problem, we demonstrate by direct numerical simulation (DNS) of the coupled fluid-PSub lattice systems sustained downstream reduction in the perturbation kinetic energy (KE), which indicates a delay in laminar-to-turbulent transition. Moreover, as dictated by the PSub lattice symmetry, this approach is immune to any changes in the direction of propagation of the instability waves and may be tuned to accommodate a relatively wide range of frequencies and wavenumbers along each direction--similar to the classical acoustic Bragg scattering problem which we also demonstrate to provide a direct analogy and comparison.
@@ Results
We form our lattice of PSubs using identical PSub units, all installed in a channel model. We select a channel flow problem for simplicity, but all the underlying concepts are readily applicable to boundary layer flows. The design of the PSub unit used in this investigation is shown in Fig. 2a. It consists of an elastic rod, composed of five unit cells, with each unit cell accommodating a local resonator (which in practice may be realized as a cantilevered beam or pillar). The dispersion curves and the amplitude and phase frequency-response characteristics for this nominal PSub configuration is shown in Fig. 2b. The product of the amplitude and phase yields a performance metric [27] which provides an a priori prediction of the behavior of the PSub once passively engaged with the flow--this quantity is also plotted in Fig. 2b.
@@ Discussion
In conclusion, the mechanism of scatterless interference reported here reveals a unique class of wave behavior that to our knowledge has not been previously reported in any branch of physics. The concept comprises subsurface phononic units arranged with full spatial flexibility and designed to vibrate at specific frequencies, enabling precise, tunable, and frequency-selective interference patterns in the overlying wave field. The physical platform we employ to demonstrate this mechanism is an unstable laminar flow field subject to perturbation waves—a canonical setting for studying aerodynamic or hydrodynamic instability. We have shown that by engineering the phonon properties of the subsurface--virtually entirely concealed beneath the flow-exposed surface--it is possible to exert full command over the very nature of the underlying fluid-structure interaction. This passive intervention offers precise mechanistic and tunable control of the behavior of wall-bounded flows in general. Furthermore, our findings demonstrate that the principles of Bloch wave analysis, traditionally applied to electronic, photonic, acoustic, and elastic waves [36, 37], can be effectively extended to flow perturbations influenced by PSubs. By leveraging locally resonant metamaterial-based individual PSub units, we have demonstrated that flow instabilities can be manipulated through interference effects, in a manner analogous to wave phenomena in periodic elastic or acoustic media. This approach enables passive control over critical processes such as the transition to turbulence, with the PSub lattice facilitating attenuation of unstable T-S waves downstream, while being robust to variations in wavenumber and direction of propagation. Other types of flow instabilities may be similarly treated in future studies. With the ongoing influx of new concepts from phonon engineering and metamaterials [38, 39, 40], the prospects of further future improvements of the performance following this framework is highly accessible, and so is the path to practical implementation given the rapid advancements in fabrication technology [41, 42]. A wide range of flow-control applications stand to benefit, spanning from hypersonic airflows to low-speed liquid transport. It is conceivable that within this framework future aircrafts will maintain a fully laminar boundary layer over their surfaces, enabling transformative improvements in fuel efficiency. In addition, PSub lattices may be simultaneously designed to inhibit aeroacoustic emissions, offering the additional rewards of improved passenger comfort and reduced environmental noise pollution. Similar advances will materialize for hydroacoustic emissions as well.